OpenStudy (thecalchater):

Fundamental Theorem of Calculus 1 questioin

OpenStudy (thecalchater):

@zepdrix @dan815

zepdrix (zepdrix):

hehe sec XD about half way through it. you're too quick for me

OpenStudy (thecalchater):

:) I think it may be b just guessing.

zepdrix (zepdrix):

Hmm I'm not getting anything close to these values... I must be doing something wrong >.<

zepdrix (zepdrix):

Oh this is the mean value theorem for INTEGRALS... not derivatives wooooops :O back to the drawing board...

OpenStudy (thecalchater):

ok XD

zepdrix (zepdrix):

Average value is obtained by\[\large\rm \frac{1}{b-a}\int\limits_a^b f(x)dx=f(c)\]Where c is somewhere between a and b.

zepdrix (zepdrix):

So umm

zepdrix (zepdrix):

\[\large\rm \frac{1}{7-1}\int\limits_1^7\frac{u^2+4}{u^2}du=?\]Understand how to do this integral? :)

OpenStudy (thecalchater):

7/396 I got.

OpenStudy (thecalchater):

any ideas?

OpenStudy (perl):

This is basically a plug into integral problem?

OpenStudy (thecalchater):

I think so.... @dan815

OpenStudy (perl):

on paper I found 11/7

OpenStudy (thecalchater):

right that is what i thought too, but I dont know if the mean value is + and - or just +

OpenStudy (perl):

We can solve f(u) = 11/7

OpenStudy (perl):

i.e. solve (u^2 + 4)/u^2 = 11/7

OpenStudy (thecalchater):

it would be both - and + sqrt7 so D?

OpenStudy (perl):

That part is tricky. The directions say that the function is defined for u between 1 and 7. Therefore I would go with just the positive.

OpenStudy (thecalchater):

ok let me submit it and see

OpenStudy (thecalchater):

I got it right thank you :)

OpenStudy (perl):

since it is talking about 'the' interval [1,7]

OpenStudy (thecalchater):

it was b only positive like I had said in the beginning... thanks for the help though

OpenStudy (perl):

:)